OpenStudy (anonymous):

Differentiate sec^2 (pi x)

OpenStudy (anonymous):

sorry it is \[y^{'}=-2{\pi}\cos^{-3}(\pi x)\sin(\pi x)\]

OpenStudy (anonymous):